OpenStudy (anonymous):

use sqrt property to solve: (5k-1)^2+126=1 k=??

OpenStudy (anonymous):

not possible squares can't be equal to negative

OpenStudy (anonymous):

I am not sure

OpenStudy (anonymous):

\[(5k-1)^2=-125\] \[5k+1=\pm\sqrt{125}=\pm5\sqrt{5}i\] \[5k=-1\pm5\sqrt{5}i\] \[k=-\frac{1}{5}\pm\sqrt{5}i\]

OpenStudy (anonymous):

ah I see complex lol sorry

OpenStudy (anonymous):

i think i saw SSG post complex questions before which is how i guessed.

OpenStudy (anonymous):

dang, wrong answer, unless I typed wrong, checking...

OpenStudy (anonymous):

the 1/5 is negative in both answers?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

hmmmm, then it seems incorrect

OpenStudy (anonymous):

nope the solution is perfect it can't be wrong

OpenStudy (anonymous):

okay, it must b my error somehow :(

OpenStudy (anonymous):

no it is postive

OpenStudy (anonymous):

i screwed up on one line and changed \[5k-1\] to \[5k+1\] by mistake. sorry

OpenStudy (anonymous):

great, no prob :)

OpenStudy (anonymous):

it is \[\frac{1}{5}-\sqrt{5}i\] and \[\frac{1}{5}+\sqrt{5}i\]

OpenStudy (anonymous):

uh oh, still not it. One more guess for me left, any idea?

OpenStudy (anonymous):

lol you taking online test

OpenStudy (anonymous):

not a test, just some online homework:)

OpenStudy (anonymous):

I'm an english student, so I can help in that area but this algebra is beyond me!!

OpenStudy (anonymous):

you could try one fraction maybe

OpenStudy (anonymous):

oh okay let me see this time I will try to keep my eyes open properly open

OpenStudy (anonymous):

\[\frac{1\pm5\sqrt{5}i}{5}\]

OpenStudy (anonymous):

it is right for sure

OpenStudy (anonymous):

it's expecting two anwers, one positive one negative. I'll retype it and see

OpenStudy (anonymous):

maybe some problem with syntax. make sure the "i" is outside of the square root

OpenStudy (anonymous):

oh crap!! that's it!!

OpenStudy (anonymous):

dang, it was right with that second form you showed me. wow, thanks for all your work!!

OpenStudy (anonymous):

whew. hope you got it right yw

OpenStudy (anonymous):

nope,lol... my bad

OpenStudy (anonymous):

:(